The transformation T : z → w in the complex plane is defined by w =
az + b z + c
, where a, b, c ∈ R.
Given that w = 3i when z = −3i, and w = 10 − 4i when z = 1 + 4i, find the values of a, b and c.
(a) Show that the points for which z is transformed to z lie on a circle and give the centre and radius of this circle.
(b) Show that the line through the point z = 4 and perpendicular to the real axis is invariant under T.
1
Expert's answer
2020-05-27T16:25:40-0400
"w={az+b\over z+c}"

Given that w = 3i when z = −3i, and w = 1 − 4i when z = 1 + 4i, find the values of a, b and c. 


"w=3i={-3ai+b\over -3i+c}""9+3ci=-3ai+b=>b=9, a=-c"

"w=1-4i={a(1+4i)+9\over 1+4i-a}""17-a+4ai=a+4ai+9""a=4, b=9, c=-4"

"w={4z+9\over z-4}"

(a) Show that the points for which "z" is transformed to "\bar{z}" lie on a circle and give the centre and radius of this circle. 


"\bar{z}={4z+9\over z-4}"

"x^2+y^2-4(x-iy)-4(x+iy)-9=0"

"x^2-8x+y^2-9=0"

"x^2-8x+16-16+y^2-9=0"

"(x-4)^2+y^2=25"

This the cartesian equation of the circle with center "(4, 0)" and radius "5."


(b) Show that the line through the point z = 4 and perpendicular to the real axis is invariant under T.

"z=4+iy"


"{4(4+iy)+9\over 4+iy-4}={25+4iy\over iy}=4-i({25\over y}), y\not=0"

Every point of the line is transformed to point of this line excepting point "z=4."